a 0.6 mm wire made of carbon and manganese can just barely support the weight of a 70.0 kg person. Suppose this wire is used to lift a 55.0 kg load. What maximum upward acceleration can be achieved without breaking the wire?
1
Expert's answer
2018-02-18T09:01:07-0500
The Newton’s second law gives
ma=T-mg
The maximum tension in the wire
T=Mg
Thus, maximum upward acceleration
a=g (M-m)/m
a=9.8 (70.0-55.0)/55.0=2.67 m/s^2